# Context-aware gate set tomography: Improving the self-consistent characterization of trapped-ion universal gate sets by leveraging non-Markovianity

### v1

@@ II Microscopic parametrization of trapped-ion GST
where, using the notation of [58] , $\tilde{\Omega}_{L,i}$ is the crossed-beam ac Stark shift for the $i$-th ion, and $\mathcal{M}_{i,m}^{z}$ is the displacement of ion $i$ in the normal vibrational mode $m$ along the $z$ axis. In this expression $\bar{z}^{0}_{m}=1/\sqrt{2m\omega_{z,m}}$ is the ground-state oscillator width, and $\eta_{\alpha,m}=\bm{\Delta k}\cdot\bm{e}_{z}\bar{z}^{0}_{m}=\sqrt{E_{R}/\omega_{z,m}}\ll 1$ is the Lamb-Dicke parameter [60].
@@
$\mathcal{J}\!\!(t_{g})=-\sum_{m}\frac{2\mathcal{F}_{i_{1},m}^{z}\mathcal{F}_{i_{2},m}^{z}}{\delta_{n}}\left(t_{g}\cos\phi^{0}_{i_{1}i_{2}}+\frac{1}{\delta_{m}}\sin(\delta_{m}t-\phi^{0}_{i_{1}i_{2}})\Big{|}_{0}^{t_{g}}\right)\!\!.$ (6)
and where we have introduced $\phi^{0}_{i_{1}i_{2}}=\Delta\bm{k}(\bm{r}_{i_{1}}^{0}-\bm{r}_{i_{2}}^{0})$. We note that one could also include counter-rotating contributions to the phonon-mediated qubit-qubit interactions, such that the time $t_{g}$ for a maximally-entangling operation $\mathcal{J}\!\!(t_{g})=\pi/4=:\theta_{LS}$ would be changed. The ideal unitary describing this evolution reads $U_{\mathrm{LS}}={\rm exp}\{-{\mathrm{i}}\theta_{LS}Z_{i_{1}}Z_{i_{2}}\}$, which generates Bell pairs from the product Hadamard basis |$\ket{\sigma=\pm}=\frac{1}{\sqrt{2}}(\ket{0}\pm\ket{1})$, namely
$U_{\mathrm{LS}}\ket{\pm,\pm\sigma}=\frac{1}{\sqrt{2}}(\ket{\pm,\pm\sigma}-{\mathrm{i}}\ket{\mp,\mp\sigma}).$ (7)
@@ II.1 Analytical channel of the noisy light-shift gate
It will prove convenient to write quantum channels on its vectorized or “natural” representation [31, 65], where density operators $\rho\in\mathcal{H}$ are written as vectors $\mathbf{vec}(\rho)=(\rho\otimes\mathbb{1}_{{\mathrm{d}im}(\mathcal{H})})\ket{\Phi}$ composed of the stacked columns (column-major order) of the matrix [66]. Here |$\ket{\Phi}=\sum_{\bm{\sigma}}\ket{\bm{\sigma}}\otimes\ket{\bm{\sigma}}/2$ with $\bm{\sigma}=(\sigma_{i_{1}},\sigma_{i_{2}})\in\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ refers to the maximally entangled state on a doubled Hilbert space $\mathcal{H}\otimes\mathcal{H}$ [10]. To analyze gate performance, we are only interested in spin dynamics and we will consequently trace over the phononic degrees of freedom. In this subspace, all factors in Eq. (5) commute. As we will discuss later, magnetic noise introduces fluctuations of the characteristic frequency of ions $\omega_{0}$, hence, the resulting channel, which we will denote as $\mathcal{E}_{\Gamma_{\mathrm{d}}}$, will also commute with Eq. (5). As detailed below, the reduced quantum channel for the complete evolution can be divided into
@@ II.1.1 Motional errors and collective dephasing
$\begin{split}U_{\mathrm{D}}={\mathrm{e}}^{\sum_{j,m}(\phi_{j,m}(t_{g})a^{\dagger}_{m}-{\rm H.c.})Z_{j}}=\sum_{\bm{\sigma}}\!\ket{\bm{\sigma}}\bra{\bm{\sigma}}\mathcal{D}(\bm{\sigma}),\end{split}$ (10)
@@
We now assume a separable initial state in which phonons follow a thermal distribution $\rho_{\mathrm{p}h}=\bigotimes_{m}\sum_{n_{m}}p_{m,n}\ket{n_{m}}\bra{n_{m}}$ |, where $p_{m,n}={\mathrm{e}}^{-\beta\omega_{m}n_{m}}/(1-{\mathrm{e}}^{\beta\omega_{m}})$ [70]. Altogether, the unitary evolution produced by $U_{\mathrm{D}}$ reads
$U_{\mathrm{D}}\rho{U_{\mathrm{D}}}^{\dagger}=\sum_{\bm{\sigma},\bm{\gamma}}\rho_{\bm{\sigma},\bm{\gamma}}\ket{\bm{\sigma}}\bra{\bm{\gamma}}\bigotimes_{m}\sum_{n_{m}}p_{m,n}\mathcal{D}(\bm{\sigma})\ket{n_{m}}\bra{n_{m}}\mathcal{D}^{\dagger}(\bm{\gamma}),$ (14)
where $\rho_{s}=\sum_{\bm{\sigma},\bm{\gamma}}\rho_{\bm{\sigma},\bm{\gamma}}\ket{\bm{\sigma}}\bra{\bm{\gamma}}$ | represents any density operator describing the qubit state of the ions. Finally, we can trace out phonons and, using the expectation value of the displacement operator for a thermal state [68], we obtain
$\mathrm{tr_{\mathrm{ph}}}\{\mathcal{E}_{\mathrm{D}}(\rho)\}=\sum_{\bm{\sigma},\bm{\gamma}}\rho_{\bm{\sigma},\bm{\gamma}}\ket{\bm{\sigma}}\bra{\bm{\gamma}}{\mathrm{e}}^{\sum_{m}\big{(}{\mathrm{i}}\Upsilon_{m}^{\bm{\sigma},\bm{\gamma}}-\mid\Phi_{m}^{\bm{\sigma},\bm{\gamma}}\mid^{2}(\bar{n}_{m}+1/2)\big{)}},$ (15)
@@
$\Gamma_{\mathrm{th}}=\left(\frac{\mid\tilde{\Omega}_{L}\mid\eta_{\alpha,m}}{\delta_{m}}\right)^{2}\left(1-\cos\delta_{m}t_{g}\right)\left(\bar{n}_{\mathrm{b}}+\textstyle\frac{1}{2}\right),$ (20)
@@ II.1.3 Effective collective Kraus channel
$\begin{split}&K_{1}=\frac{1}{2}(\sqrt{\chi_{\mathrm{d}}}+\sqrt{\chi_{\mathrm{th}}})\mathbb{1}+\frac{1}{2}(\sqrt{\chi_{\mathrm{d}}}-\sqrt{\chi_{\mathrm{th}}})Z_{i_{1}}Z_{i_{2}},\\ &K_{2}=\frac{1}{2}\sqrt{\chi_{\mathrm{th}}(1-\chi_{\mathrm{th}})}(Z_{i_{1}}-Z_{i_{2}}),\\ &K_{3}=\frac{1}{2}\sqrt{\chi_{\mathrm{d}}(1-\chi_{\mathrm{d}})}(Z_{i_{1}}+Z_{i_{2}}),\\ &K_{4}=\frac{1}{4}(1-\chi_{\mathrm{th}})(\mathbb{1}+Z_{i_{1}})(\mathbb{1}-Z_{i_{2}}),\\ &K_{5}=\frac{1}{4}(1-\chi_{\mathrm{th}})(\mathbb{1}-Z_{i_{1}})(\mathbb{1}+Z_{i_{2}}),\\ &K_{6}=\frac{1}{4}(1-\chi_{\mathrm{d}})(\mathbb{1}+Z_{i_{1}})(\mathbb{1}+Z_{i_{2}}),\\ &K_{7}=\frac{1}{4}(1-\chi_{\mathrm{d}})(\mathbb{1}-Z_{i_{1}})(\mathbb{1}-Z_{i_{2}}).\\ \end{split}$ (34)
@@ II.2 Trapped-ion two-qubit GST: A comparison of fully-general and parametrized approaches
[…]
@@
While we already presented GST during the introduction, let us further discuss its main aspects, which will be relevant for the estimation of gate set (36) with the microscopic parameterization. In a nutshell, GST provides an estimate of the gate set under study by fitting experimentally-measured frequencies to a model describing the gates. These frequencies are nothing but the outcomes of carefully designed circuits obeying a fixed structure [21]: $i)$ state preparation: the reference state $\rho_{0}$ undergoes an evolution produced by elements extracted from $\mathcal{G}$ $ii)$ base circuit: a ‘germ’, i.e., a particular combination of gates, is executed to provide sensitivity with respect to some of the parameters of the gate set. Optionally, this sequence of gates may be repeated several times to enhance precision in what it is called the long-sequence GST protocol. $iii)$ measurement: followed by some transformation such as those in state preparation, the measurement operation $M_{0}$ is applied and the results are collected.
@@
It is important to mention that stages 1)1)1 ) and 3)3)3 ) are designed to generate the informationally complete [93] sets $\big{\{}G_{s}\rho_{0}=\rho_{s}:\,s\in\mathbb{S},\,|\mathbb{S}|=d^{2}\ \big{\}}$ and $\big{\{}G_{b}M_{0}=M_{b}:\ {b}\in\mathbb{M}_{b}\big{\}}$, where the Hilbert space dimension is $d$. Each POVM $M_{b}$ has $m_{b}$ possible outcomes, such that the measurements are labeled by $\mu=(b,m_{b})\in\mathbb{M}=\mathbb{M}_{b}\times\mathbb{M}_{m_{b}}$ with $|\mathbb{M}_{b}|=d^{2}$. This means that the fiducial pairs $\big{\{}\rho_{s},M_{b}\big{\}}$ are sufficient to obtain a unique representation of any quantum channel $\hat{\mathcal{E}}$ from the probabilities $p_{\mu,s}=\mathrm{tr}\{M_{\mu}\mathcal{E}(\rho_{s})\}$. Similarly, long-sequence GST requires the set of germs to amplify every non-redundant parameter used to describe it. This condition is known as amplification completeness (AC). Once the outputs of these circuits have been collected, experimental frequencies are constructed as $f_{\gamma_{c},\mu}=N_{\gamma_{c},\mu}/N_{\gamma_{c}}$, where $N_{\gamma_{c},\mu}$ denotes the number of times that outcome $\mu$ has been obtained from circuit $\gamma_{c}$, and $N_{\gamma_{c}}$ the number of times that circuit has been executed. Finally, one estimates the gate set by fitting this data to the theoretical probabilities $p_{\gamma_{c},\mu}$ that describe those same circuits. In principle, this could be achieved by means of a simple linear inversion algorithm. Finite sampling noise, however, can lead to un-physical solutions, yielding estimates that do not satisfy the completely positive and trace preserving (CPTP) conditions of quantum maps. To solve this issue, the linear inversion estimate is often only used as an initial guess of non-convex nonlinear optimization problem, in which quantum maps can be enforced to be physical by imposing certain constraints. Different cost functions and optimization routines have been refined to circumvent local minima, and give a precise estimate at a certain resource cost [21]. More precisely, GST provides an estimate $\hat{\bm{\theta}}$ to a constrained multi-parameter point-estimation problem in which, typically, a negative log-likelihood cost function $\mathcal{C}_{\rm ML}(\bm{\theta})$ or a weighted least-squares cost function $\mathcal{C}_{\rm LS}(\bm{\theta})$ is built from the finite-frequency outcomes and then minimized
$\begin{split}&\hat{\bm{\theta}}=\texttt{argmin}\hskip 4.30554pt\begin{cases}\mathcal{C}_{\rm ML}(\bm{\theta})=-\sum\limits_{\bm{\gamma}}\sum\limits_{m_{b}}f_{\bm{\gamma}}(m_{b})\log p_{\bm{\gamma}}(m_{b}),\\ \mathcal{C}_{\rm LS}(\bm{\theta})=\sum\limits_{\bm{\gamma}}\sum\limits_{m_{b}}\frac{1}{p_{\bm{\gamma}}(m_{b})}\big{(}f_{\bm{\gamma}}(m_{b})-p_{\bm{\gamma}}(m_{b})\big{)}^{2}\!.\end{cases}\\ \end{split}$ (40)
@@ II.2.1 Scaling of long-sequence GST: single-depth estimates versus logarithmic-depth spacing
We now present the results of Figure 2, showing the scaling of precision with the maximum depth employed for long sequence (red line and data-points), using the microscopically-modelled gate set. The results follow a $m=-1$ slope tendency in log-log axis (black dashed line), consistent with the expected GST results obtained in other works [21]. As discussed in [40], and also noted in prior GST treatments [21, 32], we appreciate how this scaling saturates for sufficiently large depths, as noise begins to dominate the dynamics and we can no longer improve the estimations as the signal to noise ratio no longer improves with circuit depth.
@@
Figure 2: Average diamond distance as a function of circuit depth for the microscopic model. We run GTS on simulated data for the usual logarithmically-spaced depths GST scheme (red) and for a last-depth only scheme (blue). The number of circuits required are $N_{\mathrm{c}}=12$ for the later scheme and $N_{\mathrm{c}}=12\times\log_{2}p$ for the first scheme. We observe $1/p$ tendencies for both schemes (gray guide to the eye). We note that the noise-dominated region in which precision plateaus for the logarithmic scheme results in a degradation of precision for the last-depth only one. Parameters used for the simulations are $c=2\times 10^{9}s^{-3}$, $\tau_{c}=5\times 10^{-4}s$ and $\bar{n}_{\mathrm{b}}=5$. Each circuit is sampled $N_{\rm samples}=10000$ times. Light shaded regions represent $95\%$ confidence intervals after averaging over $100$ GST estimations for each depth.
@@ II.2.2 Sample complexity of the microscopic GST
Figure 3: Average diamond distance as a function of the total number of shots $N_{\mathrm{shots}}$ for both the microscopic model and the fully general one. We run GTS on simulated data for both models, showing a dramatic improvement in the precision of the estimates for the microscopic model at a fixed number of shots. Parameters used for the simulations are $c=2\times 10^{9}s^{-3}$, $\tau_{c}=5\times 10^{-4}s$ and $\bar{n}_{\mathrm{b}}=5$. Each circuit is sampled $N_{\rm samples}=10000$ times. Light shaded regions for the microscopic model represent $95\%$ confidence intervals after averaging over $100$ GST estimations for each depth. We complement the fully general estimates with theoretical tendency predictions for different FPR schemes, indicating the ratio of circuits employed with respect to the original design.
@@ III.1 Analytical channel of sequential light-shift gates
$\mathrm{tr_{\mathrm{ph}}}\{\mathcal{E}_{\mathrm{D}}^{(p)}(\rho)\}=\sum_{\bm{\sigma},\bm{\gamma}}\rho_{\bm{\sigma},\bm{\gamma}}\ket{\bm{\sigma}}\bra{\bm{\gamma}}\times\mathcal{A},$ (47)
@@
$\mathcal{A}=\prod_{m}\sum_{n_{m}}p_{m,n}\bra{n_{m}}\left(\mathcal{D}^{\dagger}(\bm{\gamma}){\mathrm{e}}^{{\mathrm{i}}H_{0}t_{g}}\right)^{\!\!p}\left({\mathrm{e}}^{-{\mathrm{i}}H_{0}t_{g}}\mathcal{D}(\bm{\sigma})\right)^{\!\!p}\ket{n_{m}}.$ (48)
@@
$\mathcal{A}=\prod_{m}{\mathrm{e}}^{\eta_{m}^{\bm{\sigma},\bm{\gamma}}(p)}\sum_{n_{m}}p_{m,n}\bra{n_{m}}D_{m}(\Phi^{\bm{\sigma},\bm{\gamma}}_{m}\textstyle\sum_{l=0}^{p-1}{\mathrm{e}}^{{\mathrm{i}}l\omega_{z,m}t_{g}})\ket{n_{m}},$ (49)

### v2

@@ II Microscopic parametrization of trapped-ion GST
where, using the notation of [58] , $\tilde{\Omega}_{L,i}$ is the crossed-beam ac Stark shift for the $i$-th ion, and $\mathcal{M}_{i,m}^{z}$ is the displacement of ion $i$ in the normal vibrational mode $m$ along the $z$ axis. In this expression $\bar{z}^{0}_{m}=1/\sqrt{2m\omega_{z,m}}$ is the ground-state oscillator width, and $\eta_{z,m}=\bm{\Delta k}\cdot\bm{e}_{z}\bar{z}^{0}_{m}=\sqrt{E_{R}/\omega_{z,m}}\ll 1$ is the Lamb-Dicke parameter [60].
@@
$\mathcal{J}\!\!(t_{g})=-\sum_{m}\frac{2\mathcal{F}_{i_{1},m}^{z}\mathcal{F}_{i_{2},m}^{z}}{\delta_{m}}\left(t_{g}\cos\phi^{0}_{i_{1}i_{2}}+\frac{1}{\delta_{m}}\sin(\delta_{m}t-\phi^{0}_{i_{1}i_{2}})\Big|_{0}^{t_{g}}\right)\!\!.$ (6)
[…]
@@ II.1 Analytical channel of the noisy light-shift gate
[…]
@@ II.1.1 Motional errors and collective dephasing
[…]
@@
We now assume a separable initial state in which phonons follow a thermal distribution $|\rho_{\mathrm{p}h}=\bigotimes_{m}\sum_{n_{m}}p_{m,n}\ket{n_{m}}\bra{n_{m}}$, where $p_{m,n}={\mathrm{e}}^{-\beta\omega_{m}n_{m}}/(1-{\mathrm{e}}^{\beta\omega_{m}})$ [70]. Altogether, the unitary evolution produced by $U_{\mathrm{D}}$ reads
[…]
where $|\rho_{s}=\sum_{\bm{\sigma},\bm{\gamma}}\rho_{\bm{\sigma},\bm{\gamma}}\ket{\bm{\sigma}}\bra{\bm{\gamma}}$ represents any density operator describing the qubit state of the ions. Finally, we can trace out phonons and, using the expectation value of the displacement operator for a thermal state [68], we obtain
[…]
@@
$\Gamma_{\mathrm{th}}=\left(\frac{\mid\tilde{\Omega}_{L}\mid\eta_{z,2}}{\delta_{z,2}}\right)^{2}\left(1-\cos\delta_{z,2}t_{g}\right)\left(\bar{n}_{\mathrm{2}}+\textstyle\frac{1}{2}\right),$ (20)
@@ II.1.3 Effective collective Kraus channel
After the discussion in the previous sections, we have all the necessary ingredients to compute the LS-gate quantum channel (9). The corresponding super-operator will be employed in next section to parametrize two-qubit GST, and show that one can indeed improve the precision in comparison to a fully-general approach. The super-operator representation will be convenient when dealing with the constrained optimization of GST, as the composition rule reduces to simple matrix multiplication. Let us, however, close this section by providing a Kraus decomposition [90] of the quantum channel to give further physical intuition about the noisy dynamics. We have $\rho(t_{g})=\mathcal{E}_{N}(\rho_{0})=\sum_{n}K_{n}\rho_{0}K_{n}^{\dagger}$, where $\mathcal{E}_{N}\coloneq\mathcal{E}_{N}(\Gamma_{\mathrm{d}},\Gamma_{\mathrm{th}})$ denotes the noise channel, and the Kraus operators $K_{n}$ can be expressed in terms of Pauli operators and the functions $\chi_{\mathrm{d}}={\mathrm{e}}^{-\Gamma_{\mathrm{d}}}$, $\chi_{\mathrm{th}}={\mathrm{e}}^{-\Gamma_{\mathrm{th}}}$ as follows
@@ II.2 Trapped-ion two-qubit GST: A comparison of fully-general and parametrized approaches
$\begin{split}\mathcal{G}&=\Bigl\{\rho_{0},\hskip 4.30554pt\big\{G_{LS}(\theta_{LS});\ G_{i}(\theta_{i},\phi_{i}),\,i={1,\cdots,5}\big\},\hskip 4.30554ptM_{0}\Bigr\},\end{split}$ (36)
@@
While we already presented GST during the introduction, let us further discuss its main aspects, which will be relevant for the estimation of gate set (36) with the microscopic parameterization. In a nutshell, GST provides an estimate of the gate set under study by fitting experimentally-measured frequencies to a model describing the gates. These frequencies are nothing but the outcomes of carefully designed circuits obeying a fixed structure [21]: $i)$ state preparation: the reference state $\rho_{0}$ undergoes an evolution produced by elements extracted from $\mathcal{G}$ ii)ii) base circuit: a ‘germ’, i.e., a particular combination of gates, is executed to provide sensitivity with respect to some of the parameters of the gate set. Optionally, this sequence of gates may be repeated several times to enhance precision in what it is called the long-sequence GST protocol. $iii)$ measurement: followed by some transformation such as those in state preparation, the measurement operation $M_{0}$ is applied and the results are collected.
@@
It is important to mention that stages $1)$ and $3)$ are designed to generate the informationally complete [93] sets $\big\{G_{s}\rho_{0}=\rho_{s}:\,s\in\mathbb{S},\,|\mathbb{S}|=d^{2}\ \big\}$ and $\big\{G_{b}M_{0}=M_{b}:\ {b}\in\mathbb{M}_{b}\big\}$, where the Hilbert space dimension is $d$. Each POVM $M_{b}$ has $m_{b}$ possible outcomes, such that the measurements are labeled by $\mu=(b,m_{b})\in\mathbb{M}=\mathbb{M}_{b}\times\mathbb{M}_{m_{b}}$ with $|\mathbb{M}_{b}|=d^{2}$. This means that the fiducial pairs $\big\{\rho_{s},M_{b}\big\}$ are sufficient to obtain a unique representation of any quantum channel $\hat{\mathcal{E}}$ from the probabilities $p_{\mu,s}=\mathrm{tr}\{M_{\mu}\mathcal{E}(\rho_{s})\}$. Similarly, long-sequence GST requires the set of germs to amplify every non-redundant parameter used to describe it. This condition is known as amplification completeness (AC). Once the outputs of these circuits have been collected, experimental frequencies are constructed as $f_{\gamma_{c},\mu}=N_{\gamma_{c},\mu}/N_{\gamma_{c}}$, where $N_{\gamma_{c},\mu}$ denotes the number of times that outcome $\mu$ has been obtained from circuit $\gamma_{c}$, and $N_{\gamma_{c}}$ the number of times that circuit has been executed. Finally, one estimates the gate set by fitting this data to the theoretical probabilities $p_{\gamma_{c},\mu}$ that describe those same circuits. In principle, this could be achieved by means of a simple linear inversion algorithm. Finite sampling noise, however, can lead to un-physical solutions, yielding estimates that do not satisfy the completely positive and trace preserving (CPTP) conditions of quantum maps. To solve this issue, the linear inversion estimate is often only used as an initial guess of non-convex nonlinear optimization problem, in which quantum maps can be enforced to be physical by imposing certain constraints. Different cost functions and optimization routines have been refined to circumvent local minima, and give a precise estimate at a certain resource cost [21]. More precisely, GST provides an estimate $\hat{\bm{\theta}}$ to a constrained multi-parameter point-estimation problem in which, typically, a negative log-likelihood cost function $\mathcal{C}_{\rm ML}(\bm{\theta})$ or a weighted least-squares cost function $\mathcal{C}_{\rm LS}(\bm{\theta})$ is built from the finite-frequency outcomes and then minimized
[…]
@@ II.2.1 Scaling of long-sequence GST: single-depth estimates versus logarithmic-depth spacing
We now present the results of Figure 2, showing the scaling of precision with the maximum depth employed for long sequence (red line and data-points), using the microscopically-modelled gate set. The results follow a m=1m=-1 slope tendency in log-log axis (black dashed line), consistent with the expected GST results obtained in other works [21]. As discussed in [40], and also noted in prior GST treatments [21, 32], we appreciate how this scaling saturates for sufficiently large depths, as noise begins to dominate the dynamics and we can no longer improve the estimations as the signal to noise ratio no longer improves with circuit depth.
@@
Figure 2: Average diamond distance as a function of circuit depth for the microscopic model. We run GTS on simulated data for the usual logarithmically-spaced depths GST scheme (red) and for a last-depth only scheme (blue). The number of circuits required are $N_{\mathrm{c}}=12$ for the later scheme and $N_{\mathrm{c}}=12\times\log_{2}p$ for the first scheme. We observe $1/p$ tendencies for both schemes (gray guide to the eye). We note that the noise-dominated region in which precision plateaus for the logarithmic scheme results in a degradation of precision for the last-depth only one. Parameters used for the simulations are $c=2\times 10^{9}s^{-3}$, $\tau_{c}=5\times 10^{-4}s$ and $\bar{n}_{\mathrm{b}}=5$. Each circuit is sampled $N_{\rm samples}=10000$ times. Light shaded regions represent $95\%$ confidence intervals after averaging over 100100 GST estimations for each depth.
@@ II.2.2 Sample complexity of the microscopic GST
Figure 3: Average diamond distance as a function of the total number of shots $N_{\mathrm{shots}}$ for both the microscopic model and the fully general one. We run GTS on simulated data for both models, showing a dramatic improvement in the precision of the estimates for the microscopic model at a fixed number of shots. Parameters used for the simulations are $c=2\times 10^{9}s^{-3}$, $\tau_{c}=5\times 10^{-4}s$ and $\bar{n}_{\mathrm{b}}=5$. Each circuit is sampled $N_{\rm samples}=10000$ times. Light shaded regions for the microscopic model represent $95\%$ confidence intervals after averaging over 100100 GST estimations for each depth. We complement the fully general estimates with theoretical tendency predictions for different FPR schemes, indicating the ratio of circuits employed with respect to the original design.
@@ III.1 Analytical channel of sequential light-shift gates
[…]
@@
[…]
@@
[…]